How do i find the slope of the line with equation 9y=6, showing my work ?
Found 2 solutions by umnikmo, solver91311:
Answer by umnikmo(75)

this line is y=6/9=2/3
this line is parallel to x axis, therefore slope =0

Answer by solver91311(24713)



In , the equation , which is the simplified form of your given equation simply describes the set of all ordered pairs where the -coordinates are all equal to and the -coordinates are just the set of Real numbers. Since the coordinates are all equal, no matter which pair of ordered pairs you choose for calculating the slope, the numerator of the slope fraction will be zero, hence the slope is zero.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism